OpenStudy (anonymous):

Would someone help me with this equation? I just started learning how to do them, and am a bit confused... x+2 over x+3 divided by 3x over 4x.

OpenStudy (anonymous):

\[\frac{ x+2}{ x+3}\div \frac{ 3x }{ 4x }\]

OpenStudy (anonymous):

x+2 is x2, x+3 is x3, which leaves you with x2/x3 divided by 3x/4x

OpenStudy (anonymous):

Alright, then what do I do?

OpenStudy (anonymous):

\[\frac{ 2x }{3x}\div \frac{ 3x }{ 4x }\]

OpenStudy (anonymous):

first change 3x/4x to .......4x/3x this makes it a multiplcation then multiply the top by 4x and the bottom by 3x

OpenStudy (anonymous):